OpenStudy (supersquidsisters):

what is the square root of 1/2 as a fraction??

OpenStudy (lord_box):

\[\sqrt{1/2} = \frac{ \sqrt{1} }{ \sqrt{2} } = \frac{ 1 }{ \sqrt{2}}\] Rationalize this.

OpenStudy (supersquidsisters):

@Lord_Box would you leave the denominator as a decimal?

OpenStudy (lord_box):

No, we typically don't leave roots on the bottom. This is an example of rationalization:\[\frac{ 1 }{ \sqrt(5)} \times \frac{ \sqrt(5)}{ \sqrt(5) } = \frac{\sqrt(5)}{5}\]

OpenStudy (supersquidsisters):

@Lord_Box oh, so it would be square root of 2 over 2?

OpenStudy (lord_box):

That's right

OpenStudy (supersquidsisters):

oh, ok!! thank you so much!!

OpenStudy (lord_box):

You're welcome